# Optimal Transportation by Orthogonal Coupling Dynamics

### v1

@@ 2.1. Main Dynamics
the marginals of $p_{t}$ remain unchanged. However, the general form of this tangent space, given by $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$ in Eq. (7), does not provide us with a concise and convenient projection (this avenue was explored, to some degree, in [4]). Yet, if instead of $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$, we focus on a sub-manifold, which still preserves the marginals, the problem can become tractable. In particular, the tangent space $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)=(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}$ offers a straight-forward projection, while preserving the marginals. Noting that the Riemannian gradient is the orthogonal projection of the Euclidean one onto the tangent space [15], we get the projected version of $v^{\textrm{gd}}$, as shown in Fig. 1.
Figure 1. Geometric illustration of OCD; (a) The gradient descent, which leaves the polytope $\Pi(\mu,\nu)$, is projected onto the tangent space $(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}$; (b) Initial condition as product Gaussian probability density; (c) Evolved by $v^{\textrm{ocd}}$, the initial density is updated by projected gradient descent in $L^{2}$-cost; (d) Evolved by $v^{\textrm{gd}}$, the initial density is updated by gradient descent in $L^{2}$-cost, leaving $\Pi(\mu,\nu)$.
@@
which we refer to as the Orthogonal Coupling Dynamics (OCD). The orthogonality comes from the observation that $\mathbb{E}[v_{1}^{\textrm{ocd}}\cdot X_{t}]=\mathbb{E}[v_{2}^{\textrm{ocd}}\cdot Y_{t}]=0$. In general, there is no guarantee that OCD reaches global minimum of (1), as the projection onto the sub-manifold $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ can prevent the dynamics from visiting all possible couplings between $\mu$ and $\nu$. This can be readily noticed, e.g. if $Y_{0}$ is initialized as a sub-optimal map of $X_{0}$. As a result, the dynamics remains frozen at a sub-optimal fixed point. However, as we will discuss, there are several nuances in (2.1), which warrant investigation of OCD on both theoretical and numerical fronts. In particular, we observe that OCD ensures the cost is non-increasing. This result plays a crucial role in two important properties that we deduce: we observe that sub-optimal couplings turn out to be unstable (see Proposition 2.3), and that OCD gives the sharpest decay in $L^{2}$-cost among processes with velocity in $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ (see Theorem 2.6). In practice, the dynamics reaches accurate estimations of the optimization problem (1), as we demonstrate in results section.
From conceptual point of view, by replacing the optimization problem (1) with the OCD system (2.1), we turn the Monge-Kantorovich problem into a series of regression problems, implied by the conditional expectation. While the original problem, given by Eq. (1), is a constrained optimization over distributions, the OCD entails an unconstrained optimization over functions, in $L^{2}$. This transformation allows us to leverage tools of conditional expectation estimation in order to achieve efficient computation of optimal transport in contrast to the original constrained infinite dimensional linear programming problem.
However, before analyzing OCD, let us mention that alternatively, one can consider the Langevin processes
@@
where $\epsilon>0$ is the regularization parameter; $W_{t},B_{t}$ are independent Wiener processes with values in $\mathbb{R}^{n}$; and $f_{\mu,\nu}$ are densities of the measures $\mu,\nu$. Assuming log-concave densities, Conforti et al [23] have shown that the measure generated by the Langevin dynamics (2.1) converges to the solution of (1), once the latter is regularized by entropy with parameter $\epsilon$. Despite the theoretical grounding of SDE (2.1), we have three main motivations behind focusing on OCD (2.1), instead:
@@
The open question remains under which technical assumptions the case of $\epsilon=0$, which is OCD (2.1), would converge to the solution of the Monge-Kantorovich problem. While we do not directly address this problem and despite this knowledge gap, we present new results on both theoretical and numerical aspects of OCD, presenting its basic features along with computational relevance for practical optimal transport problems.
@@ 2.2. General Structural Properties
The OCD dynamics (2.1) leads to a certain favourable theoretical features. These are rooted in the fact that the conditional expectation is a projection operator in $L^{2}$ with self-adjoint and contraction characteristics. We leverage these characteristics to derive theoretical properties of the OCD. While the detailed justifications are deferred to Supplementary Information, we provide an overview of the main theoretical results below, where the proofs are provided in Supplementary Information.
@@
    Marginal Preservation. Since the velocities $v_{1,2}^{\textrm{ocd}}$, introduced in Eq. (2.1), are orthogonal with respect to the functions of $X_{t}$ and $Y_{t}$, respectively, a direct computation shows that the distributions of $X_{t}$ and $Y_{t}$ remain unchanged. More precisely we have the following result.
@@
    with velocities $v=(v_{1},v_{2})\in\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$, and the initial condition $(X_{0},Y_{0})\sim\pi$, where $\pi\in\Pi(\mu,\nu)$ and $p_{t}$ is their joint law. Therefore $p_{t}\in\Pi(\mu,\nu)$.
@@
    Consider $(X_{t},Y_{t})$ to be the solution of OCD (2.1) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. We have
@@
    In other words, the OCD (2.1) monotonically decreases the cost $\mathbb{E}[c(X_{t},Y_{t})]$ until the latter reaches its stationary value. This property is crucial to show that the sub-optimal fixed points of the dynamics are unstable (as discussed in the following).
@@
    $\Pi^{*}(\mu,\nu)\mathrel{\mathop{:}}=\left\{\pi\in\Pi(\mu,\nu)\Big{|}\int_{\mathbb{R}^{2n}}c(x,y)\pi(dx,dy)>\int_{\mathbb{R}^{2n}}c(x,y)\pi^{\textrm{opt}}(dx,dy)\right\}$ (17)
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (2.1) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. Suppose that the joint measure of $(X_{t},Y_{t})$ fulfills
@@
    at some $T>0$. The coupling $p_{T}$ is unstable in the sense that if the law is perturbed by
@@
    where $10<\epsilon<$ 1, the solution $(X_{t},Y_{t})$ would not have the measure $p_{T}$ for any time later.
@@
    McKean-Vlasov System. The OCD (2.1) can be equivalently interpreted as a probability flow in the sense of Eq. (5). Due to the absence of Brownian motion, this probability flow is in the form of the Vlasov equation. The latter arises in the mean field limit of interacting particles, e.g. see [33, 49]. By exploiting the projection operator $\mathsf{S}$, the corresponding probability density evolution takes the concise form
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (2.1) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. Suppose $p_{t}$ is the joint law of $(X_{t},Y_{t})$ with the density $\rho_{t}$. Therefore $\rho_{t}$ is weak solution of the Vlasov-type equation Eq. (20).
    We expect that the Vlasov equation (20) admits a variational formulation in terms of Otto calculus [35, 3, 57, 9], if appropriate set of couplings is considered (those that can be generated by $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$). However, we leave this interesting question to a separate work. Instead, we will present a variational formulation of OCD in the probability space for the $L^{2}$ Monge-Kantorovich cost, in the next section.
@@ 2.3. Key Features in $L^{2}$ Monge-Kantorovich
So far, we discussed that OCD (2.1) brings a given joint distribution closer to the solution of the Monge-Kantorovich problem. In other words, the dynamics decreases the cost $\mathbb{E}[c(X_{t},Y_{t})]$. In this section, we present a finer result on the decay of the quadratic cost $c(x,y)=\lVert x-y\rVert_{2}^{2}$, resulted from OCD, i.e. for the ODE system
@@
In this scenario, the optimal transport problem is equivalent to maximizing correlation between $X_{t}$ and $Y_{t}$, see [52, 54]. Focusing on (22), we show the following results.
@@
[…]
@@
[…]
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (2.1) with the initial condition $(X_{0},Y_{0})\sim\mu\otimes\nu$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. Therefore
@@
    Sharp Descent and Variational Formulation. The $L^{2}$-OCD (22) gives rise to a sharp decay in the cost $\mathbb{E}[c(X_{t},Y_{t})]$. More precisely, if we restrict the set of velocities to the tangent space $(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}$, the OCD gives fastest decay in the cost $\mathbb{E}[c(X_{t},Y_{t})]$.
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (15) and the initial condition $(X_{0},Y_{0})\in\pi$, where $\pi\in\Pi(\mu,\nu)$. Let
[…]
@@
[…]
@@
[…]

### v2

@@ 2.1. Main Dynamics
the marginals of $p_{t}$ remain unchanged. However, the general form of this tangent space, given by $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$ in Eq. (1.3), does not provide us with a concise and convenient projection (this avenue was explored, to some degree, in [angenent2003minimizing]). Yet, if instead of $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$, we focus on a sub-manifold, which still preserves the marginals, the problem can become tractable. In particular, the tangent space $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)=\ker(\mathsf{P}_{p_{t},X})\times\ker(\mathsf{P}_{p_{t},Y})$ offers a straight-forward projection, while preserving the marginals. Noting that the Riemannian gradient is the orthogonal projection of the Euclidean one onto the tangent space [boumal2023introduction], we get the projected version of $v^{\textrm{gd}}$, as shown in Fig. 1.
Figure 1. Geometric illustration of OCD; (a) The gradient descent, which leaves the polytope $\Pi(\mu,\nu)$, is projected onto the tangent space $\ker(\mathsf{P}_{p_{t},X})\times\ker(\mathsf{P}_{p_{t},Y})$; (b) Initial condition as product Gaussian probability density; (c) Evolved by $v^{\textrm{ocd}}$, the initial density is updated by projected gradient descent in $L^{2}$-cost; (d) Evolved by $v^{\textrm{gd}}$, the initial density is updated by gradient descent in $L^{2}$-cost, leaving $\Pi(\mu,\nu)$.
@@
which we refer to as the Orthogonal Coupling Dynamics (OCD). The orthogonality comes from the observation that $\mathbb{E}[v_{1}^{\textrm{ocd}}\cdot X_{t}]=\mathbb{E}[v_{2}^{\textrm{ocd}}\cdot Y_{t}]=0$. In general, there is no guarantee that OCD reaches global minimum of (2), as the projection onto the sub-manifold $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ can prevent the dynamics from visiting all possible couplings between $\mu$ and $\nu$. This can be readily noticed, e.g. if $Y_{0}$ is initialized as a sub-optimal map of $X_{0}$. As a result, the dynamics remains frozen at a sub-optimal fixed point. However, as we will discuss, there are several nuances in (14), which warrant investigation of OCD on both theoretical and numerical fronts. In particular, we observe that OCD ensures the cost is non-increasing. This result plays a crucial role in two important properties that we deduce: we observe that sub-optimal couplings turn out to be unstable (see Proposition 2.3), and that OCD gives the sharpest decay in $L^{2}$-cost among processes with velocity in $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ (see Theorem 2.6). In practice, the dynamics reaches accurate estimations of the optimization problem (2), as we demonstrate in results section.
From conceptual point of view, by replacing the optimization problem (2) with the OCD system (14), we turn the Monge-Kantorovich problem into a series of regression problems, implied by the conditional expectation. While the original problem, given by Eq. (2), is a constrained optimization over distributions, the OCD entails an unconstrained optimization over functions, in $L^{2}$. This transformation allows us to leverage tools of conditional expectation estimation in order to achieve efficient computation of optimal transport in contrast to the original constrained infinite dimensional linear programming problem.
However, before analyzing OCD, let us mention that alternatively, one can consider the Langevin dynamics
@@
where $\epsilon>0$ is the regularization parameter; $W_{t},B_{t}$ are independent Wiener processes with values in $\mathbb{R}^{n}$; and $f_{\mu,\nu}$ are densities of the measures $\mu,\nu$. Assuming log-concave densities, Conforti et al [conforti2023projected] have shown that the measure generated by the Langevin dynamics (2.1) converges to the solution of (2), once the latter is regularized by entropy with parameter $\epsilon$. Despite the theoretical grounding of SDE (2.1), we have three main motivations behind focusing on OCD (14), instead:
@@
The open question remains under which technical assumptions the case of $\epsilon=0$, which is OCD (14), would converge to the solution of the Monge-Kantorovich problem. While we do not directly address this problem and despite this knowledge gap, we present new results on both theoretical and numerical aspects of OCD, presenting its basic features along with computational relevance for practical optimal transport problems.
@@ 2.2. General Structural Properties
The OCD dynamics (14) leads to a certain favourable theoretical features. These are rooted in the fact that the conditional expectation is a projection operator in $L^{2}$ with self-adjoint and contraction characteristics. We leverage these characteristics to derive theoretical properties of the OCD. We provide an overview of the main theoretical results below, where the proofs are given in section˜3.
@@
    Marginal Preservation. Since the velocities $v_{1,2}^{\textrm{ocd}}$, introduced in Eq. (14), are orthogonal with respect to the functions of $X_{t}$ and $Y_{t}$, respectively, a direct computation shows that the distributions of $X_{t}$ and $Y_{t}$ remain unchanged. More precisely we have the following result.
@@
    with velocities $v=(v_{1},v_{2})\in\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$, and the initial condition $(X_{0},Y_{0})\sim\pi$, where $\pi\in\Pi(\mu,\nu)$ and $p_{t}$ is their joint law. Under the regularity assumptions stated in Assumption 3.1, $p_{t}\in\Pi(\mu,\nu)$.
@@
    Consider $(X_{t},Y_{t})$ to be the solution of OCD (14) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$. Under the regularity assumptions stated in Assumption 3.1, we have
@@
    In other words, the OCD (14) monotonically reduces the cost $\mathbb{E}[c(X_{t},Y_{t})]$ until the latter reaches its stationary value. This property is crucial to show that the sub-optimal fixed points of the dynamics are unstable (as discussed in the following).
@@
    $\Pi^{*}(\mu,\nu)\mathrel{\mathop{:}}=\left\{\pi\in\Pi(\mu,\nu):\int_{\mathbb{R}^{2n}}c(x,y)\pi(dx,dy)>\int_{\mathbb{R}^{2n}}c(x,y)\pi^{\textrm{opt}}(dx,dy)\right\}$
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (14) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$. Suppose that the joint measure of $(X_{t},Y_{t})$ fulfills
@@
    at some $T>0$. Under the regularity assumptions stated in Assumption 3.1, the coupling $p_{T}$ is unstable in the sense that if the law is perturbed by
@@
    where $0<\epsilon<1$, the solution $(X_{t},Y_{t})$ would not have the measure $p_{T}$ for any time later.
@@
    McKean-Vlasov System. The OCD (14) can be equivalently interpreted as a probability flow in the sense of Eq. (6). Due to the absence of Brownian motion, this probability flow is in the form of the Vlasov equation. The latter arises in the mean field limit of interacting particles, e.g. see [golse2003mean, paul2022microscopic]. By exploiting the projection operator $\mathsf{S}$, the corresponding probability density evolution takes the concise form
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (14) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$. Suppose $p_{t}$ is the joint law of $(X_{t},Y_{t})$ with the density $\rho_{t}$. Under the regularity assumptions stated in Assumption 3.1, $\rho_{t}$ is weak solution of the Vlasov-type equation Eq. (20).
    We expect that the Vlasov equation (20) admits a variational formulation in terms of Otto calculus [jordan1998variational, ambrosio2007gradient, santambrogio2017euclidean, bashiri2020gradient], if appropriate set of couplings is considered (those that can be generated by $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$). However, we leave this interesting question to a separate work. Instead, we will present a variational formulation of OCD in the probability space for the $L^{2}$ Monge-Kantorovich cost, in the next section.
@@ 2.3. Key Features in $L^{2}$ Monge-Kantorovich
So far, we discussed that OCD (14) brings a given joint distribution closer to the solution of the Monge-Kantorovich problem. In other words, the dynamics decreases the cost $\mathbb{E}[c(X_{t},Y_{t})]$. In this section, we present a finer result on the decay of the quadratic cost $c(x,y)=\lVert x-y\rVert_{2}^{2}$, resulted from OCD, i.e. for the ODE system
@@
In this scenario, the optimal transport problem is equivalent to maximizing correlation between $X_{t}$ and $Y_{t}$, see [rachev2006mass, ruschendorf1990characterization]. Focusing on (22), we show the following results.
@@
    $S^{n}_{+}$ $\mathrel{\mathop{:}}=$ $\left\{A\in\mathbb{R}^{n\times n}|A=A^{T},A\succeq 0\right\}\ .$ (23)
@@
    $J_{t}$ $\mathrel{\mathop{:}}=$ $\mathbb{E}\left[X_{t}^{\prime}\otimes{Y_{t}^{\prime}}\right]$ (24)
@@
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (14) with the initial condition $(X_{0},Y_{0})\sim\mu\otimes\nu$, given the regularity assumptions stated in Assumption 3.1. Therefore
@@
    Sharp Descent and Variational Formulation. The $L^{2}$-OCD (22) gives rise to a sharp decay in the cost $\mathbb{E}[c(X_{t},Y_{t})]$. More precisely, if we restrict the set of velocities to the tangent space $\ker(\mathsf{P}_{p_{t},X})\times\ker(\mathsf{P}_{p_{t},Y})$, the OCD gives fastest decay in the cost $\mathbb{E}[c(X_{t},Y_{t})]$.
@@
    Under the regularity assumptions stated in Assumption 3.1, consider $(X_{t},Y_{t})$ to be the solution of the ODE system (16) and the initial condition $(X_{0},Y_{0})\in\pi$, where $\pi\in\Pi(\mu,\nu)$. Let
    $\mathcal{C}(p_{t})$ $\vcentcolon$= $\frac{1}{2}\int_{\mathbb{R}^{2n}}\lVert x-y\rVert_{2}^{2}p_{t}(dx,dy)$ (26)
@@
    $\mathcal{L}_{p_{t}}(v)$ $\vcentcolon$= $\frac{d}{dt}\mathcal{C}(p_{t})+\frac{1}{2}\int_{\mathbb{R}^{2n}}\lVert v\rVert_{2}^{2}p_{t}(dx,dy)\ ,$ (27)
@@
    $v^{opt}_{p_{t}}$ $\mathrel{\mathop{:}}=$ $\operatorname*{arg\,min}_{v\in\ker(\mathsf{P}_{p_{t},X})\times\ker(\mathsf{P}_{p_{t},Y})}\mathcal{L}_{p_{t}}(v)\ .$ (28)